OpenStudy (anonymous):

Solve the equation below for x. 4x-5y=20

OpenStudy (saifoo.khan):

What basically you have to do is to bring x on one side of the equation. Agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh i got it

OpenStudy (anonymous):

wait but i still need a little help so hold on a minute

OpenStudy (anonymous):

after y = -4 - 4x/5 what next

OpenStudy (anonymous):

wait no i am solving for x not y

OpenStudy (anonymous):

i guess first i have to find y right

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

you can try to use http://www.wolframalpha.com if that is useful

OpenStudy (anonymous):

it is not very useful

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

i think you cant solve an equation with two unknowns u will need as many linear equations as possible

OpenStudy (anonymous):

ok the answer choices r

OpenStudy (anonymous):

F. x = 5/4y+5

OpenStudy (anonymous):

Gx = 4/5y+25

OpenStudy (anonymous):

sorry that was G. = to

OpenStudy (anonymous):

sorry again that was G. x = to

OpenStudy (anonymous):

ok i get it now we are looking for x so why make it more complicated what you do is

OpenStudy (anonymous):

4x - 5y=20 so from there you add 5 to both sides +5y +5y then you do that and that equals 4x = 20 + 5y

OpenStudy (anonymous):

then from there you divide 4 on all 3 sides 4x = 20 + 5 - - - 4 4 4

OpenStudy (anonymous):

4 cancels out and you are left with x = 20/4 + 5/4 then you divide 20/4 that equals 5 and then you have

OpenStudy (anonymous):

4x = 5 + 5/4 and that is ur answer actually it is 4x=5/4+5

OpenStudy (anonymous):

sorry it equals x= 5/4y + 5

OpenStudy (anonymous):

and that is your answer

OpenStudy (anonymous):

any questions r comments

OpenStudy (anonymous):

so the answer is F. x = 5/4y+5

OpenStudy (anonymous):

y dont u accept messages @skullpatrol